The diameter of 20th dark ring in a Newton's Ring system viewed normally is 0.6 c.m. Calculate the thickness of corresponding air film if wavelength of light used is 6X10 cm.


1
Expert's answer
2021-01-22T12:56:34-0500

Let's write the condition for dark ring:


"D_n=\sqrt{4n\lambda R}."

From this formula we can find "R":


"R=\dfrac{D_n^2}{4n\lambda}=\dfrac{(6.0\cdot10^{-3}\ m)^2}{4\cdot20\cdot6\cdot10^{-7}\ m}=0.75\ m."

Finally, we can find the thickness of corresponding air film:


"t=\dfrac{r_n^2}{2R}=\dfrac{(3.0\cdot10^{-3}\ m)^2}{2\cdot 0.75\ m}=6\cdot10^{-6}\ m=6\ \mu m."

Answer:

"t=6\ \mu m."